a body is projected vertically up with a velocity 60m/s, the % of its initial ke converted into pe after 3 seconds?

Grade:11

1 Answers

Vikas TU
7 years ago
at 3 seconds , h = 60*3 – 0.5g*9 = > 135 meter.
P.E. = mgh = >1350m
K.E = 0.5m*3600 = > 1800m
% = 1350m/1800m * 100 => 75%